OpenStudy (anonymous):

if f(x) = cos2x, find f(3pi/4).

OpenStudy (anonymous):

try to use these three trigo identities related to cos(2x) 1. \[\cos ^{2}(x)-\sin ^{2}(x)\] 2. \[2\cos ^{2}(x)-1\] 3. \[1-2\sin ^{2}(x)\] place \[x= (3\pi/4)\] and you will get the same answer for all three of the equation.

OpenStudy (anonymous):

yeah, n remember to set your calculator to radian mode.

OpenStudy (anonymous):

thank you. :)